A student, starting from rest, slides down a water slide. On the way down, a kinetic frictional force (a nonconservative force) acts on her. The student has a mass of 81.0 kg, and the height of the water slide is 10.9 m. If the kinetic frictional force does -7.70 × 103 J of work, how fast is the student going at the bottom of the slide?
Answer #1

Potential energy, U = mgh = 81 x 9.8 x 10.9 = 8652.42 J

Change in energy, dE = 8652.42 - 7700 = 952.42 J

0.5 x 81 x v^2 = 952.42

Velocity, v = 4.85 m/s

Comment in case any doubt please rate my answer...